# PrE-Text: Training Language Models on Private Federated Data in the Age of LLMs

### v1

@@ 1 Introduction
(1) PrE-Text (Private Evolution-Text) algorithm. We propose PrE-Text, a new algorithm for DP synthetic text generation. We build on the following insights: (1) PE must generate variations of samples (e.g., similar images). We adapt this requirement to the language domain by carefully utilizing mask-filling models (Devlin et al., 2018; Lewis et al., 2019) instead of the diffusion models they used for images. (2) PE alone does not generate enough high-quality synthetic data to effectively finetune an LLM. Hence, we add a post-processing phase, in which we use the outputs of PE to seed high-quality LLMs trained on public data to generate more similar text. [^1] In concurrent work, Xie et al. (2024) extend PE to the text domain using similar techniques to ours, developing a new algorithm called Aug-PE. Some differences: (1) Their focus is purely on generating better DP synthetic text data, whereas ours is on understanding whether DP synthetic data can take the place of on-device learning. To this end, they study the centralized setting, while we study the federated setting. (2) PrE-Text uses the original PE as a subroutine in a larger algorithm, whereas Aug-PE is Private Evolution with core components rewritten. (3) Their approach relies on a closed-source model (ChatGPT) for consistent improvements over baselines, whereas ours uses open-source models (LLaMA). This matters because prompts for synthetic data generation that work on ChatGPT did not work well on LLaMA, and the OpenAI API policy might not allow training on outputs (OpenAI, 2024). Done carefully, this can generate orders of magnitude more useful training data, greatly aiding generalization (all without incurring more privacy cost, due to the post-processing property of DP (Dwork, 2006)).
@@ 4.1 Models stored on-device
Expand-private $\epsilon=\infty$ 0.730 $\pm$ 0.000 0.684 $\pm$ 0.000 0.689 $\pm$ 0.000 0.687 $\pm$ 0.000
@@
1.920 $\pm$ 0.001
Expand-private $\epsilon=\infty$ 1.374 $\pm$ 0.002 1.682 $\pm$ 0.001 1.668 $\pm$ 0.001 1.585 $\pm$ 0.001
@@
Non-finetuned $\epsilon=0$ 0.510 $\pm$ 0.000 0.466 $\pm$ 0.000 0.473 $\pm$ 0.000 0.460 $\pm$ 0.000
PrE-Text $\epsilon=1.29$ 0.522 $\pm$ 0.000 0.479 $\pm$ 0.001 0.484 $\pm$ 0.000 0.473 $\pm$ 0.001
PrE-Text $\epsilon=7.58$ 0.523 $\pm$ 0.000 0.480 $\pm$ 0.000 0.483 $\pm$ 0.000 0.473 $\pm$ 0.000
Expand-private $\epsilon=\infty$ 0.526 $\pm$ 0.000 0.484 $\pm$ 0.000 0.488 $\pm$ 0.001 0.476 $\pm$ 0.000
@@
Non-finetuned $\epsilon=0$ 2.915 $\pm$ 0.000 3.250 $\pm$ 0.000 3.220 $\pm$ 0.000 3.218 $\pm$ 0.000
PrE-Text $\epsilon=1.29$ 2.811 $\pm$ 0.000 3.177 $\pm$ 0.004 3.133 $\pm$ 0.002 3.110 $\pm$ 0.000
PrE-Text $\epsilon=7.58$ 2.795 $\pm$ 0.008 3.164 $\pm$ 0.005 3.139 $\pm$ 0.008 3.101 $\pm$ 0.000
Expand-private $\epsilon=\infty$ 2.746 $\pm$ 0.002 3.085 $\pm$ 0.001 3.088 $\pm$ 0.001 3.026 $\pm$ 0.000
@@ Appendix A Additional Related Work
(Private) federated learning. Hou et al. (2021b) study how to do saddle point optimization in the federated setting. Weller et al. (2022) study the multilingual language model learning problem in federated learning. Ramaswamy et al. (2020) train production next-token-prediction models with differential privacy. Xu et al. (2023a) study the private federated learning setting, modifying private on-device training to be more noise and communication efficient. Gupta et al. (2022) show an attack on how to recover text used to train federated text models (if differential privacy is not used). Liu et al. (2023) study how to obtain low-error estimates of the aggregated histogram across user histograms under differential privacy.
@@ Appendix C Dataset generation details
We also evaluate our method on a question-and-answer dataset focused on coding and technical topics, which has text content partitioned by users. We construct 1250 users by making a federated dataset where each client has the comments associated with that user in the dataset. For each user, we cap the number of comments to 128. This forms the training dataset. The evaluation dataset is made up of the first 2000 samples of the next 100 users from the same dataset. We call this dataset Code.

### v3

@@ 1 Introduction
(1) PrE-Text (Private Evolution-Text) algorithm. We propose PrE-Text, a new algorithm for DP synthetic text generation. We build on the following insights: (1) PE must generate variations of samples (e.g., similar images). We adapt this requirement to the language domain by carefully utilizing mask-filling models (Devlin et al., 2018; Lewis et al., 2019) instead of the diffusion models they used for images. (2) PE alone does not generate enough high-quality synthetic data to effectively finetune an LLM. Hence, we add a post-processing phase, in which we use the outputs of PE to seed high-quality LLMs trained on public data to generate more similar text. [^1] In concurrent work, Xie et al. (2024) extend PE to the text domain using similar techniques to ours, developing a new algorithm called Aug-PE. Some differences: (1) Their focus is on generating better DP synthetic text data, whereas ours is on understanding whether DP synthetic data can take the place of on-device learning. To this end, they study the centralized setting, while we study the federated setting. (2) PrE-Text uses the original PE as a subroutine in a larger algorithm, whereas Aug-PE is Private Evolution with core components rewritten. Done carefully, this can generate orders of magnitude more useful training data, greatly aiding generalization (all without incurring more privacy cost, due to the post-processing property of DP (Dwork, 2006)).
@@ 4.1 Models stored on-device
Expand-private $\epsilon=\infty$ 0.730 $\pm$ 0.000 0.684 $\pm$ 0.000 0.689 $\pm$ 0.000 0.680 $\pm$ 0.000
@@
2.027 $\pm$ 0.000
1.927 $\pm$ 0.002
Expand-private $\epsilon=\infty$ 1.374 $\pm$ 0.002 1.682 $\pm$ 0.001 1.668 $\pm$ 0.001 1.677 $\pm$ 0.001
@@
Non-finetuned $\epsilon=0$ 0.458 $\pm$ 0.000 0.466 $\pm$ 0.000 0.473 $\pm$ 0.000 0.460 $\pm$ 0.000
PrE-Text $\epsilon=1.29$ 0.522 $\pm$ 0.000 0.479 $\pm$ 0.001 0.484 $\pm$ 0.000 0.475 $\pm$ 0.000
PrE-Text $\epsilon=7.58$ 0.523 $\pm$ 0.000 0.480 $\pm$ 0.000 0.483 $\pm$ 0.000 0.476 $\pm$ 0.000
Expand-private $\epsilon=\infty$ 0.526 $\pm$ 0.000 0.484 $\pm$ 0.000 0.488 $\pm$ 0.001 0.479 $\pm$ 0.000
@@
Non-finetuned $\epsilon=0$ 3.252$\pm$ 0.000 3.250 $\pm$ 0.000 3.220 $\pm$ 0.000 3.218 $\pm$ 0.000
PrE-Text $\epsilon=1.29$ 2.811 $\pm$ 0.000 3.177 $\pm$ 0.004 3.133 $\pm$ 0.002 3.145 $\pm$ 0.000
PrE-Text $\epsilon=7.58$ 2.795 $\pm$ 0.008 3.164 $\pm$ 0.005 3.139 $\pm$ 0.008 3.135 $\pm$ 0.000
Expand-private $\epsilon=\infty$ 2.746 $\pm$ 0.002 3.085 $\pm$ 0.001 3.088 $\pm$ 0.001 3.054 $\pm$ 0.000
@@ Appendix A Additional Related Work
On-device training. Hou et al. (2021b) study how to do saddle point optimization in the federated setting. Weller et al. (2022) study the multilingual language model learning problem in federated learning. Ramaswamy et al. (2020) train production next-token-prediction models with differential privacy. Xu et al. (2023a) study the private federated learning setting, modifying private on-device training to be more noise and communication efficient. Gupta et al. (2022) show an attack on how to recover text used to train federated text models (if differential privacy is not used).
@@ Appendix C Dataset generation details
We also evaluate our method on a question-and-answer dataset focused on coding and technical topics, which has text content partitioned by users. We construct 1250 users by making a federated dataset where each client has the comments associated with that user in the dataset. For each user, we cap the number of comments to 128. This forms the training dataset. The evaluation dataset is made up of the first 2000 samples of the next 100 users from the same dataset. We call this dataset Code. The results on the Code evaluation dataset differ slightly from the version published in ICML 2024, due to an earlier error in the train/test split, which has been corrected. The overall conclusions remain the same.